The probability that the randomly selected student gets a GED is _______

GED: In a certain high school, the probability that a student drops out is 0.05, and the probability that a dropout gets a high school equivalency diploma (GED) is 0.15. What is the probability that a randomly selected student gets a GED? Round your answer to four decimal places, if necessary.


The probability that the randomly selected student gets a GED is _______ 
Answer #1

Given that in a certain high school, the probability that a student drops out is 0.05, and the probability that a dropout gets a high school equivalency diploma (GED) is 0.15. We have to find the probability that a randomly selected student gets a GED.

Thus, here, Probability that a student drops out, P(drops out) = 0.05. -------(1)

Now, the probability that the student gets GED given that the student is a dropout, P(gets GED|drops out) = 0.15 . -----------(2)

Thus, from the multiplication rule, we get,

Probability that a student gets a GED = P(drops out) * P(gets GED|drops out) = 0.05*0.15 = 0.0075 . [From (1) and (2)]

Thus, the probability that a randomly selected student gets a GED = 0.0075 .